### v1

@@ Abstract
Automating architectural floorplan design is vital for housing and interior design, offering a faster, cost-effective alternative to manual sketches by architects. However, existing methods, including rule-based and learning-based approaches, face challenges in design complexity and constrained generation with extensive post-processing, and tend to obvious geometric inconsistencies such as misalignment, overlap, and gaps. In this work, we propose a novel generative framework for vector floorplan design via structural graph generation, called GSDiff, focusing on wall junction generation and wall segment prediction to capture both geometric and semantic aspects of structural graphs. To improve the geometric rationality of generated structural graphs, we propose two innovative geometry enhancement methods. In wall junction generation, we propose a novel alignment loss function to improve geometric consistency. In wall segment prediction, we propose a random self-supervision method to enhance the model’s perception of the overall geometric structure, thereby promoting the generation of reasonable geometric structures. Employing the diffusion model and the Transformer model, as well as the geometry enhancement strategies, our framework can generate wall junctions, wall segments and room polygons with structural and semantic information, resulting in structural graphs that accurately represent floorplans. Extensive experiments show that the proposed method surpasses existing techniques, enabling free generation and constrained generation, marking a shift towards structure generation in architectural design.
@@ Introduction
We represent the vector floorplan as a structural graph (Figure 1 (a)): wall junctions are abstracted as nodes, and wall segments as edges. $G=(V,E)$, where $V=\{v_{1},\ldots,v_{n}\}$ is the node set, and $E=\{(v_{i},v_{j})|v_{i},v_{j}\in V\}$ is the edge set. $v_{i}=(c_{i},r_{i},b_{i})$, where $c_{i}=(x_{i},y_{i})\in\mathbb{R}^{2}$ is the positional coordinate in the range of $[-1,1)$. $r_{i}\in[0,1]^{n}$ denotes the semantics of all rooms that surround $v_{i}$, where 0 for absence and 1 for presence of the specific room category, and $n$ is the number of room categories, which is 7 (Living room, Bedroom, Kitchen, Bathroom, Balcony, Storage, External area) in our experiments. Let $F=\{R_{1},R_{2},\ldots,R_{m}\}$ represent all polygonal loops of rooms, each loop $R_{i}=\{v_{i,1},v_{i,2},\ldots,v_{i,N_{i}}|v_{i,j}\in V\}$ is defined by a sequence of graph nodes. The room category of $R_{i}$ is the shared room category of all graph nodes forming that room. Different from (Sun et al. 2022), we set a fixed node number $N$ by considering background nodes. Background nodes are filtered out by assigning a background attribute for each node, denoted as $b_{i}\in\{0,1\}$, where 0 denotes the junction node and 1 for the background node. Vector floorplans have a multi-level structure. Spatially, 0D wall junctions form 1D wall segments, which close into 2D rooms. Semantically, each wall junction that forms a room has the same room semantics. Therefore, generating all 0D wall junctions basically determines the whole floorplan structure.
@@
$Alg(\hat{V}_{0})=\sum_{i=1}^{n}\min\left(g(\Delta b^{X}_{i}),g(\Delta b^{Y}_{i})\right)$ (4)
$Alg^{k}(\hat{V}_{0})=\sum_{i=1}^{n}\sum_{j=1}^{s}\left(Base^{k}\left(\min\left(g^{k}(\Delta b^{X}_{i}),g^{k}(\Delta b^{Y}_{i})\right)\right)\right)_{i}$ (5)
@@
Table 1: Quantitative evaluation. Each experiment is repeated 5 times to eliminate randomness, and the average results are reported. The smaller the better for all metrics. The best results are highlighted in bold. Void cells: GED is only applicable to topological graph constraints.
Constraint Method FID KID GED Living Kitchen Bedroom Storage Bathroom Balcony
Topology House-GAN++ 48.40 54.66 3.9 0.860 1.300 0.984 1.151 1.378 1.037
HouseDiffusion 11.87 7.23 2.59 0.955 0.971 0.979 0.567 0.967 0.994
Ours 6.82 1.79 0.49 0.992 0.991 1.004 0.752 0.966 0.976
Boundary Graph2Plan 8.40 1.34 - 1.034 0.975 1.013 0.727 0.959 0.859
WallPlan 9.07 1.02 - 0.992 1.033 0.923 1.136 1.058 1.008
Ours 7.50 0.56 - 1.007 0.990 0.992 0.454 0.967 0.935

### v2

@@ Abstract
Automating architectural floorplan design is vital for housing and interior design, offering a faster, cost-effective alternative to manual sketches by architects. However, existing methods, including rule-based and learning-based approaches, face challenges in design complexity and constrained generation with extensive post-processing, and tend to obvious geometric inconsistencies such as misalignment, overlap, and gaps. In this work, we propose a novel generative framework for vector floorplan design via structural graph generation, called GSDiff, focusing on wall junction generation and wall segment prediction to capture both geometric and semantic aspects of structural graphs. To improve the geometric rationality of generated structural graphs, we propose two innovative geometry enhancement methods. In wall junction generation, we propose a novel alignment loss function to improve geometric consistency. In wall segment prediction, we propose a random self-supervision method to enhance the model’s perception of the overall geometric structure, thereby promoting the generation of reasonable geometric structures. Employing the diffusion model and the Transformer model, as well as the geometry enhancement strategies, our framework can generate wall junctions, wall segments and room polygons with structural and semantic information, resulting in structural graphs that accurately represent floorplans. Extensive experiments show that the proposed method surpasses existing techniques, enabling free generation and constrained generation, marking a shift towards structure generation in architectural design. Code and data are available at https://github.com/SizheHu/GSDiff.
@@ Introduction
We represent the vector floorplan as a structural graph (Figure 1 (a)): wall junctions are abstracted as nodes, and wall segments as edges. $G=(V,E)$, where $V=\{v_{1},\ldots,v_{n}\}$ is the node set, and $E=\{(v_{i},v_{j})|v_{i},v_{j}\in V\}$ is the edge set. $v_{i}=(c_{i},r_{i},b_{i})$, where $c_{i}=(x_{i},y_{i})\in\mathbb{R}^{2}$ is the positional coordinate in the range of $[-1,1)$. $r_{i}\in[0,1]^{n}$ denotes the semantics of all rooms that surround $v_{i}$, where 00 for absence and $1$ for presence of the specific room category, and $n$ is the number of room categories, which is $7$ (Living room, Bedroom, Kitchen, Bathroom, Balcony, Storage, External area) in our experiments. Let $F=\{R_{1},R_{2},\ldots,R_{m}\}$ represent all polygonal loops of rooms, each loop $R_{i}=\{v_{i,1},v_{i,2},\ldots,v_{i,N_{i}}|v_{i,j}\in V\}$ is defined by a sequence of graph nodes. The room category of $R_{i}$ is the shared room category of all graph nodes forming that room. Different from (Sun et al. 2022), we set a fixed node number $N=53$ by considering background nodes, which is determined by the maximum number of nodes in floorplans in our dataset. Background nodes are filtered out by assigning a background attribute for each node, denoted as $b_{i}\in\{0,1\}$, where 0 denotes the junction node and 1 for the background node. Vector floorplans have a multi-level structure. Spatially, 0D wall junctions form 1D wall segments, which close into 2D rooms. Semantically, each wall junction that forms a room has the same room semantics. Therefore, generating all 0D wall junctions basically determines the whole floorplan structure.
@@
$\text{Alg}(\hat{V}_{0})=\sum_{i=1}^{n}g(\min\left(\Delta b_{i}^{X},\Delta b_{i}^{Y}\right))$ (4)
$Alg^{2}(\hat{V}_{0})=\sum_{i=1}^{n}g^{2}(\sum_{j=1}^{s}\left(Base^{2}\left(\min\left(\Delta b^{X}_{i},\Delta b^{Y}_{i}\right)\right)\right)_{j})$ (5)
@@
Table 1: Quantitative evaluation. # Param: parameter counts. Each experiment is repeated 5 times to eliminate randomness, and the average results are reported. The smaller the better for all metrics. The color cyan and magenda mark the top-two results. Void cells: GED is only applicable to topological graph constraints.
Constraint Method # Param FID KID GED Living Kitchen Bedroom Storage Bathroom Balcony
Topology House-GAN++ 2M 48.40 54.66 3.9 0.860 1.300 0.984 1.151 1.378 1.037
HouseDiffusion 27M 11.87 7.23 2.59 0.955 0.971 0.979 0.567 0.967 0.994
Ours (27M) 27M 6.64 1.62 0.57 0.977 0.981 0.989 0.817 0.954 0.966
Ours 125M 6.82 1.79 0.49 0.992 0.991 1.004 0.752 0.966 0.976
Boundary Graph2Plan 8M 8.40 1.34 - 1.034 0.975 1.013 0.727 0.959 0.859
WallPlan 106M 9.07 1.02 - 0.992 1.033 0.923 1.136 1.058 1.008
Ours (106M) 106M 7.83 0.51 - 1.000 0.954 1.007 0.500 0.964 0.977
Ours 137M 7.50 0.56 - 1.007 0.990 0.992 0.454 0.967 0.935
@@
Retrieval Analysis
To further evaluate the generalization ability of the model, we performed a comprehensive retrieval analysis of the generated floorplans against those in the training set. For each generated floorplan (3,000 in total), we compute the minimum Euclidean distance and Wasserstein distance in image space between it and all floorplans in the training set: if the training set sample corresponding to the minimum distance is visually similar to the generated floorplan, it indicates possible overfitting. The retrieval results show that there is a significant difference between the results we generate and those retrieved from the dataset, i.e., our model has good generalization performance. However, the distance values are not intuitive enough. To illustrate the difference intuitively, we provide several examples in Fig.  7 showing the closest match between the generated floorplans and the ones in the training set.
Figure 7: Retrieval Analysis. Generated: the generated floorplans. Euclidean(Wasserstein): the closest training floorplans based on Euclidean(Wasserstein) distance.
Discussion on Model Sizes
To ensure a fair comparison, we evaluated our method under equivalent model parameter sizes.
Topology-constrained generation
Our original model for topology-constrained generation has approximately 125 million parameters, while House-GAN++ and HouseDiffusion have about 2 million and 27 million parameters, respectively. To assess the impact of model size, we reduced our model’s parameter count to 27 million to match that of HouseDiffusion, retraining it with all other configurations unchanged. As shown in Figure 6 and Table 1, even with the same parameter size, our method still achieves superior performance. This indicates that the effectiveness of our method stems from its inherent advantages rather than a larger parameter count. Regarding House-GAN++, we could not adjust our model to match their 2 million parameters. However, House-GAN++ has inherent problems. See the supplementary material for details.
Boundary Constrained Generation
Our model has about 137 million parameters. In comparison, WallPlan has about 106 million parameters and Graph2Plan has about 8 million parameters. To evaluate the impact of model size, we reduce the number of parameters of our model to 106 million to match WallPlan and retrain it. As shown in Fig. 5 and Table 1, our method still achieves superior performance, which also shows that our method has its inherent advantages. Regarding Graph2Plan, we were unable to adjust the model to match their 8 million parameters. However, Graph2Plan has inherent problems. See the supplementary material for details.
Evaluation on Other Datasets
To further demonstrate the generalization and effectiveness of our model, we conducted experiments on the LIFULL dataset (LIFULL Co. 2016). We obtained the vectorized subset of LIFULL dataset via Raster-to-Graph (Hu et al. 2024), which contains 10,804 floorplans: 500 are used for validation, 500 for testing, and the rest for training. We trained our unconstrained model using the same configurations as with RPLAN, adjusting only the room categories to align with the 12 in LIFULL. Table 3 and Fig. 8 showcases the quantitative and qualitative results, respectively, demonstrating our method’s generalizability and effectiveness in diverse scenarios. However, as the Raster-to-Graph authors themselves mentioned, the annotation quality of the LIFULL vectorized dataset they extracted is weaker and the number is much smaller than that of RPLAN, which make the effect on LIFULL limited. Possible future works would include extending our method to the MSD (van Engelenburg et al. 2024) dataset, which contains a significant share of layouts of multi-apartment dwellings. MSD lacks vectorized structural graphs, and the presence of complex shapes of walls in the images makes vectorized structural graphs difficult to extract.
Table 3: Quantitative evaluation on the LIFULL dataset.
Dataset FID KID FE(%) AE
LIFULL 12.44 3.61 6.11 3.75
Figure 8: Qualitative evaluation on the LIFULL dataset.
@@ Conclusion
Acknowledgments
We would like to thank the anonymous reviewers for their constructive suggestions and comments. This work is supported by the National Natural Science Foundation of China (62102126, 62372153) and the Fundamental Research Funds for the Central Universities of China (JZ2023HGTB0269, JZ2022HGQA0163). In this paper, we used ”LIFULL HOME’S Dataset” provided by LIFULL Co., Ltd. via IDR Dataset Service of National Institute of Informatics.